Solve the problem involving speed of transverse wave in string with complete solutions. A guitar string has a length of .75 m and a mass of 4.8 g when subjected to a tension of 350 N. Calculate for the speed of the wave.


1
Expert's answer
2021-12-24T13:48:12-0500

We can find the velocity of a transverse wave in a string from the formula:


"v=\sqrt{\dfrac{T}{\mu}}=\sqrt{\dfrac{TL}{m}},""v=\sqrt{\dfrac{350\ N\times0.75\ m}{4.8\times10^{-3}\ kg}}=234\ \dfrac{m}{s}."